Problem 5-8
Yield to Maturity and Call with Semiannual Payments

Thatcher Corporation's bonds will mature in 15 years. The bonds have a face value of $1,000 and an 7.5% coupon rate, paid semiannually. The price of the bonds is $900. The bonds are callable in 5 years at a call price of $1,050. Round your answers to two decimal places.

What is their yield to maturity?
%

What is their yield to call?
%
Answer #1
Yield to maturity Yield to call
Par value of bonds 1000 Par value of bonds 1000
Issue of bonds 900 Issue of bonds 900
Semi annuall interest 37.5 Semi annuall interest 37.5
Number of period to maturity 30 call price 1050
Number of periods to call 10 Number of periods to call 10
Annuity PVF for 30 periods at 4.35% 16.58029 Annuity PVF for 10 periods at 5.46% 7.5521
PVF for 30th period at 4.35% 0.278757 PVF for 10th period at 5.46% 0.587655
Present value of interest 621.7609 Present value of interest 283.2038
Present value of maturity 278.757 Present value of maturity 617.0378
Price of bonds 900.5179 Price of bonds 900.2415
Annual Yield to maturity = 4.35*2 = 8.70% Annual Yield to call = 5.46*2=10.92%